I'm drowning in Logs & Inverses: My question states:
Determine if this is a one-to-one. If yes, find the formula for the inverse.
f(x) = (5x+4)/(7x+6)
I have listed it as a one-to-one and have written the following:
Changed f(x) to "y"
y =(5x+4)/(7x+6)
Then substituted "x" for "y" to get
x = (5y+4)/(7y+6)
My problem is that I don't know how to write the inverse.
Please help.

Answer by scott8148(6628)

solve for y

multiplying by 7y+6 __ 7xy+6x=5y+4 __ subtracting 6x+5y __ 7xy-5y=4-6x

factoring __ y(7x-5)=4-6x __ dividing by 7x-5 __ y=(4-6x)/(7x-5) __ this is the inverse